OpenStudy (anonymous):

Use the graph of its related function to write the solutions of each inequality. -x2 +10x -25> 0

OpenStudy (anonymous):

Rewrite the inequality as x^2 - 10x + 25 < 0 The expression x^2 - 10x + 25 = (x - 5)^2 The graph of y = (x - 5)^2 is a parabola y = x^2 moved 5 units to the right, that is the vertex would be at (5,0). Now, you look at that parabola, and detrmine for what values of x is that parabola greater than 0.

OpenStudy (anonymous):

Correction in last statement..look and detrmine for what values of x is the parabola less than 0